# MapExRL: Human-Inspired Indoor Exploration with Predicted Environment Context and Reinforcement Learning

### v1

@@ Abstract
Path planning for robotic exploration is challenging, requiring reasoning over unknown spaces and anticipating future observations. Efficient exploration requires selecting budget-constrained paths that maximize information gain. Despite advances in autonomous exploration, existing algorithms still fall short of human performance, particularly in structured environments where predictive cues exist but are underutilized. Guided by insights from our user study, we introduce MapExRL, which improves robot exploration efficiency in structured indoor environments by enabling longer-horizon planning through reinforcement learning (RL) and global map predictions. Unlike many RL-based exploration methods that use motion primitives as the action space, our approach leverages frontiers for more efficient model learning and longer horizon reasoning. Our framework generates global map predictions from the observed map, which our policy utilizes, along with the prediction uncertainty, estimated sensor coverage, frontier distance, and remaining distance budget, to assess the strategic long-term value of frontiers. By leveraging multiple frontier scoring methods and additional context, our policy makes more informed decisions at each stage of the exploration. We evaluate our framework on a real-world indoor map dataset, achieving up to an 18.8% improvement over the strongest state-of-the-art baseline, with even greater gains compared to conventional frontier-based algorithms.
@@ II-B Map Prediction for Exploration
These approaches allow robots to reason beyond observed areas and focus on exploring areas guided by predictions; however, they still struggle due to their their action selection strategy, particularly when handling environments of varying sizes and topological complexities or operating under budget constraints. In this work, we aim to address these limitations.
@@ IV-B Quantitative Results
TABLE I: Table 1: Results of the user study compared to MapEx [7]. Values reflect the average final reward (higher is better) and the 95% confidence interval.
@@ V Approach
Figure 4: The MapExRL pipeline, where observed maps are processed through three independent global map prediction models, generating prediction maps. These maps are averaged and passed through a convolutional encoder to extract a 256-dimensional feature vector. This vector is concatenated with frontier centers, prediction and utility scores, distances from the agent, and remaining budget. The resulting vector is fed into a fully connected network that outputs $N$ values, and the argmax is selected as the index of the frontier action.
@@ V-A Map Encoder
The 128×128 pixel map is processed by an encoder with three convolutional layers and a fully connected layer, producing a $C$=256 dimensional latent vector. Separate map encoders are used for the actor and critic.
@@ V-C Reward Function
[…]
@@ V-F Baselines
Our first baseline is the conventional frontier-based exploration method [1], where the agent chooses the nearest frontier based on Euclidean distance, without utilizing any global map predictions. Our second baseline, MapEx [7], selects the frontier with the highest information gain, calculated by jointly considering sensor coverage and variance from predicted maps, normalized by the Euclidean distance. We also include a random exploration baseline, where the agent randomly selects a frontier from the top $N$ frontiers in the RL observation space. This baseline validates that our model learns to navigate efficiently using the top $N$ frontiers rather than randomly picking from the list.
@@ VI Results
Map 2 is an irregular L-shaped building that is challenging to explore, creating numerous opportunities for backtracking, which is a common issue with the baseline methods. In contrast, our policy often effectively selected paths that minimized backtracking and efficiently explored hard-to-predict areas using the global map prediction models. A sample trajectory across all methods on Map 2 is illustrated in Fig. 5, demonstrating the efficiency of our policy in topologically challenging environments. As observed, the global map prediction models struggled to predict the bottom-left part of the building accurately, and our RL model was the only method to successfully explore that area.

### v2

@@ Abstract
Path planning for robotic exploration is challenging, requiring reasoning over unknown spaces and anticipating future observations. Efficient exploration requires selecting budget-constrained paths that maximize information gain. Despite advances in autonomous exploration, existing algorithms still fall short of human performance, particularly in structured environments where predictive cues exist but are underutilized. Guided by insights from our user study, we introduce MapExRL, which improves robot exploration efficiency in structured indoor environments by enabling longer-horizon planning through a learned policy and global map predictions. Unlike many learning-based exploration methods that use motion primitives as the action space, our approach leverages frontiers for more efficient model learning and longer horizon reasoning. Our framework generates global map predictions from the observed map, which our policy utilizes, along with the prediction uncertainty, estimated sensor coverage, frontier distance, and remaining distance budget, to assess the strategic long-term value of frontiers. By leveraging multiple frontier scoring methods and additional context, our policy makes more informed decisions at each stage of the exploration. We evaluate our framework on a real-world indoor map dataset, achieving up to an 18.8% improvement over the strongest state-of-the-art baseline, with even greater gains compared to conventional frontier-based algorithms. Website: mapexrl.github.io
@@ II-B Map Prediction for Exploration
These approaches allow robots to reason beyond observed areas and focus on exploring areas guided by predictions; however, they still struggle due to their action selection strategy, particularly when handling environments of varying sizes and topological complexities or operating under budget constraints. In this work, we aim to address these limitations.
@@ IV-B Quantitative Results
TABLE I: Results of the user study compared to MapEx [7]. Values reflect the average final reward (higher is better) and the 95% confidence interval.
@@ V Approach
Figure 4: The MapExRL pipeline, where observed maps are processed through three independent global map prediction models, generating prediction maps. These maps are averaged and passed through a convolutional encoder to extract a 256-dimensional feature vector. This vector is concatenated with frontier centers, prediction and utility scores, distances from the agent, and remaining budget. The resulting vector is fed into a fully connected network that outputs NN values, and the argmax is selected as the index of the frontier action.
@@ V-A Map Encoder
The 128×128 pixel map is processed by an encoder with three convolutional layers and a fully connected layer, producing a CC=256 dimensional latent vector. Separate map encoders are used for the actor and critic.
@@ V-C Reward Function
We use the Intersection over Union (IoU) of the occupied cells for the global map prediction as our metric of performance. IoU is computed as $\text{IoU}=\frac{\text{TP}}{\text{FP}+\text{TN}+\text{TP}}$, where $\text{TP},\text{FP},\text{TN}$ denote true positives, false positives, and true negatives, respectively. To avoid penalizing minor map discrepancies—like wall thickness variations, gaps, dataset artifacts, or small unobservable spaces—we use binary dilations as a buffer in IoU calculation.
@@ V-F Baselines
Our first baseline is the conventional frontier-based exploration method [1], where the agent chooses the nearest frontier based on Euclidean distance, without utilizing any global map predictions. Our second baseline, MapEx [7], selects the frontier with the highest information gain, calculated by jointly considering sensor coverage and variance from predicted maps, normalized by the Euclidean distance. We also include a random exploration baseline, where the agent randomly selects a frontier from the top $N$ frontiers in the observation space. This baseline validates that our model learns to navigate efficiently using the top $N$ frontiers rather than randomly picking from the list.
@@ VI Results
Map 2 is an irregular L-shaped building that is challenging to explore, creating numerous opportunities for backtracking, which is a common issue with the baseline methods. In contrast, our policy often effectively selected paths that minimized backtracking and efficiently explored hard-to-predict areas using the global map prediction models. A sample trajectory across all methods on Map 2 is illustrated in Fig. 5, demonstrating the efficiency of our policy in topologically challenging environments. As observed, the global map prediction models struggled to predict the bottom-left part of the building accurately, and our approach was the only method to successfully explore that area.